Part A 45g of gasoline are mixed with 45g of water. What is the average density...


Part A


45g of gasoline are mixed with 45g of water. What is the average density of the mixture?


Express your answer using two significant figures.




Part B


55cm3 of gasoline are mixed with 60cm3 of water. What is the average density of the mixture?


Express your answer using two significant figures.
Answer #1

part A



the density of gasoline is approx 0.75g/cm^3 (it varies depending on the exact type of gas); the density of water is 1g/cm^3

the weighted average of densities is


avg density = (45g x 0.75g/cm^3 + 45g x1 g/cm^3)/(45g+45g) = 0.875 g/cm^3




part B


55g of gasoline are mixed with 60 g of water. What is the average density of the mixture.
first u need the gasoline and water density from tables:

gasoline density = 0.7g/cm3
water density = 1g/cm3


water volume = 60g /(1g/cm3) = 60cm3
gasoline volume = 55 / (0.7g/cm3) = 78.57cm3

total volume = 60 + 78.57 = 138.57 cm3

total mass = 55 + 60 = 115g

density of mix = 115 /138.57 = 0.829 kg/m3